.Discrete math. Any idea?

If you want a poll to be accurate within 5 percent with 90 percent confidence, what roughly is the number of people we should ask.
Answer #1

Number of people (n) is given by:

n=\frac{Z^{2}_{\alpha /2}\times p\times (1-p)}{e^{2}}

Given:

\alpha = 0.10

From Table critical values of Z = \pm 1.645

e = 0.05

Since p is not given, take p = 0.5.

Substituting, we get:

n=\frac{1.645^{2}\times 0.5\times (1-0.5)}{0.05^{2}}=271

So,

Answer is:

271